The resistivity of a potentiometer wire is if the area of cross section of the wire is 4 cm square the current flowing in a circuit is one ampere the potential gradient is 7.5 volt per metre

Grade:12

1 Answers

Khimraj
5 years ago
 
Potential gradient = p*I/A
So 7.5 = p*1/4*10-4
so p = 3*10-3 ohm*m
Hope it clears.....................